### v1

# Perturbations of Einstein-Maxwell-phantom spacetime: Instabilities of charged Ellis wormholes and quasinormal modes of black holes

Abstract
Phantom scalar fields, as a viable candidate for dark energy, have been instrumental in eliminating spacetime singularities and constructing wormholes and regular black holes. We investigate the Einstein-Maxwell-phantom (EMP) framework, in which the Ellis wormholes can be charged and regular black holes can be admitted. While the previous study has shown the stability of EMP wormholes under massless scalar field perturbations, we further perform a comprehensive linear analysis of the EMP spacetime through phantom scalar field perturbations and gravito-electromagnetic field perturbations. Our analyses of effective potentials and finite difference time profiles reveal the linear instability of EMP wormholes. In the black hole scenario, the quasinormal spectra of the EMP theory, where the matrix-valued direct integration method and the Prony method are used, recover those of general relativity when the scalar charge goes to zero. By introducing the concepts of generalized specific charge and mixing angle, we quantify how the relative contributions between the phantom scalar and the electromagnetic fields produce distinct modifications in the quasinormal spectra.
1 Introduction
Ever since the spectroscopic and photometric observations of Type Ia supernovae in the late 20th century, the cosmological data have revealed [1, 2, 3, 4, 5] that the universe is undergoing accelerated expansion. The acceleration is typically attributed to dark energy, whose equation of state is defined by the ratio of its pressure $p$ and energy density $\rho$, i.e.,
$w\equiv p/\rho,$ (1)
with $w<-1/3$ required for cosmic acceleration. Within the widely accepted $\Lambda$CDM model, dark energy is realized as the cosmological constant $\Lambda$, for which $w=-1$ [6]. Alternative scenarios, such as quintessence models [7, 8] and K-essence models [9, 10], allow $w>-1$, whereas phantom dark energy $(w<-1)$ would lead to a future Rip [11, 12, 13, 14, 15]. Although the current observations show remarkable consistency with $\Lambda$CDM, they do not exclude a slightly phantom behavior: several independent probes suggest $w\sim-1.03$ [16, 17, 18], and a combined analysis yields $w=-1.013^{+0.038}_{-0.043}$ [19].
In Einstein’s general theory of relativity, the acceleration of the universe can be accounted for either by adding a negative cosmological constant $\Lambda$ on the left-hand side of the Einstein equation or by introducing exotic matter with negative pressure to the source term on the right-hand side [12]. The former prescription realizes Gliner’s $\mu$-vacuum [20] in a de Sitter geometry, which can prevent spacetime singularity arising in gravitational collapse. A spacetime singularity, where the matter density and spacetime curvature diverge, marks the breakdown of physics. According to the Penrose singularity theorem [21], a singularity is inevitable within the framework of general relativity and under reasonable energy conditions, but no naked singularity exists, i.e., a singularity is hidden by event horizons. If the latter prescription is adopted, source terms that violate certain energy conditions are indispensable in order to construct regular (nonsingular) black holes. Notable examples include magnetically charged solutions in nonlinear electrodynamics [22, 23, 24] and configurations with phantom scalar hairs [25, 26, 27, 28, 29] (see Ref. [30] for a brief review).
The existence of black holes has been empirically confirmed through the historical direct detection of gravitational waves from a binary black hole merger by Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo [31], with numerous subsequent events reinforcing this statement [32, 33, 34, 35]. Moreover, Fermi Gamma-ray Burst Monitor (Fermi-GBM) and International Gamma-Ray Astrophysics Laboratory (INTEGRAL) observed [37, 36] a short gamma-ray burst from the same event in 1.74 seconds after LIGO and Virgo detected gravitational waves from the binary neutron star merger — GW170817. Furthermore, the first images of the supermassive black hole at the center of M87 [38, 39, 40, 41, 42, 43] and Sagittarius A* in the Milky Way [44, 45, 46, 47, 48, 49], captured by Event Horizon Telescope (EHT), inaugurate an era of multi-messenger astronomy. These breakthroughs provide us with more possibilities to test fundamental physical problems, e.g., the singularity problem [50, 51] and the validity of the no-hair theorem [52, 53, 54].
Black holes, as nearly out-of-equilibrium dissipative systems [55], resonate in a series of quasinormal modes (QNMs) when perturbed or formed after coalescence, a stage also known as ringdown [56, 57]. Theoretical study and experimental observations of black hole ringdown is crucial for probing gravity in the strong field regime, testing general relativity and constraining modified gravity theories [58, 51, 59, 60]. Tentative echoes observed in LIGO/Virgo data from binary black hole mergers have been reported[61], sparking widespread discussion [63, 64, 62, 65, 66]. Echo proposals often invoke quantum modifications near an event horizon [67, 68, 73, 69, 70, 71, 72], driven by the fact that classical black holes, which lack a natural reflection mechanism, do not inherently produce echo signals, while wormholes naturally generate echoes during the ringdown stage [74, 75, 76, 78, 77].
As with regular black holes, wormhole models also require exotic matter that violates energy conditions, specifically the weak energy condition [80, 81]. The first eponymous wormhole solution was developed by Ellis [82] in a theory where the gravity is coupled to a massless phantom scalar field. In general, a phantom field is characterized by a flapping sign in the kinetic term of its action [83]. More recently, a charged Ellis wormhole was proposed [84] by the way to extend a class of Einstein-Maxwell-dilaton (EMD) theory [85] to an Einstein-Maxwell-phantom (EMP) theory through a complex transformation. In this EMP spacetime, sufficient electric charge yields a wormhole geometry, while below a critical charge the solution becomes a regular black hole, which can be classified as Type I or Type II based on its horizon structure. Further study [78] indicates that, under massless scalar field perturbations, EMP wormholes remain stable and exhibit echo signals near the wormhole-black hole transition threshold. Note that the EMP spacetime is sourced by the phantom scalar field and the Maxwell field, and the latter can also be phantom-like in certain region. Such phantom-like behaviors imply negative kinetic energy of the source terms and can potentially induce spacetime instabilities. Investigating both phantom scalar and gravito-electromagnetic field perturbations is therefore essential for assessing the dynamical stability of the EMP spacetime and may offer fresh insights into gravity.
In this work, we focus on the axial (gravito-electromagnetic) field perturbation of the EMP spacetime. The organization of this paper is as follows. In Sec. 2, we briefly review the derivation and properties of the EMP spacetime. In Sec. 3, we study the linear perturbations of the EMP spacetime and derive the master equations of phantom scalar field perturbations and gravito-electromagnetic field perturbations. In Sec. 4, we investigate the instability of EMP wormholes. In Sec. 5, we compute the quasinormal frequencies (QNFs) of Type I black holes using the matrix-valued direct integration method and the Prony method. Finally, the conclusion is made in Sec. 6.
2 Einstein-Maxwell-phantom spacetime
The EMP theory can be obtained from the EMD theory, where the latter arises [79] as a four-dimensional low-energy effective action of string theory,
$S_{\mathrm{EMD}}=\int\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}^{4}x\sqrt{-g}\left(R-2\left(\partial\phi_{\mathrm{d}}\right)^{2}-\frac{1}{4}\mathrm{e}^{-2\phi_{\mathrm{d}}}F^{2}\right),$ (2)
where $g$ is the metric determinant, $R$ is the Ricci scalar and the dilaton scalar field $\phi_{\mathrm{d}}$ is minimally coupled to gravity, but nonminimally coupled to the Maxwell field $F=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}A$ in the form $\mathrm{e}^{-2\phi_{\mathrm{d}}}$. Rescaling the dilaton $\phi=2\phi_{\mathrm{d}}$ yields the standard coupling $\mathrm{e}^{-\phi}$. Moreover, such a nonminimal coupling is further generalized [85] to $\left(\gamma_{1}\cosh{\phi}+\gamma_{2}\sinh{\phi}\right)^{-1}$, where $\gamma_{1}$ and $\gamma_{2}$ are coupling constants, and the generalized coupling function reduces to $\mathrm{e}^{-\phi}$ when $\gamma_{1}=\gamma_{2}=1$. Under the following complex transformation,
$\left(\phi,\gamma_{2}\right)\to\left(\mathrm{i}\phi,\mathrm{i}\gamma_{2}\right),$ (3)
one arrives [84] at the EMP action,
$S=\int\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}^{4}x\sqrt{-g}\left(R+\frac{1}{2}(\partial\phi)^{2}-\frac{1}{4}Z^{-1}F^{2}\right),\qquad Z=\gamma_{1}\cos\phi+\gamma_{2}\sin\phi,$ (4)
where the scalar field is phantom-like due to its flapping sign of the kinetic term. In addition, since the coupling function $Z$ is not positive definite regardless the value of $\gamma_{1}$ and $\gamma_{2}$, the Maxwell field acquires a phantom-like behavior in negative regions of $Z$. The equations of motion associated with the variation of the phantom scalar field $\phi$, the Maxwell field $A_{\mu}$ and the gravitational field $g_{\mu\nu}$ are given respectively by
$\square\phi=\frac{1}{4}\frac{\partial Z^{-1}}{\partial\phi}F^{2},$ (5)
$\nabla_{\mu}(Z^{-1}F^{\mu\nu})=0,$ (6)
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=T_{\mu\nu}^{A}+T_{\mu\nu}^{\phi},$ (7)
where the energy-momentum tensors of the Maxwell field $T_{\mu\nu}^{A}$ and the phantom scalar field $T_{\mu\nu}^{\phi}$ take the forms,
$T_{\mu\nu}^{A}=\frac{1}{2}Z^{-1}\left(F_{\mu\nu}^{2}-\frac{1}{4}g_{\mu\nu}F^{2}\right),\qquad T_{\mu\nu}^{\phi}=-\frac{1}{2}\partial_{\mu}\phi\partial_{\nu}\phi+\frac{1}{4}g_{\mu\nu}(\partial\phi)^{2}.$ (8)
Assuming a static and spherically symmetric wormhole metric,
$\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}s^{2}$ $=-h(r)\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}t^{2}+\frac{1}{h(r)}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r^{2}+(r^{2}+q^{2})\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\Omega^{2}_{2},$ (9)
one gives [84] the charged Ellis wormhole solution of the EMP theory,
$\phi=2\arccos\left(\frac{r}{\sqrt{r^{2}+q^{2}}}\right),\qquad F_{tr}=-\frac{QZ}{r^{2}+q^{2}},\qquad h(r)=1-\frac{\gamma_{2}Q^{2}r}{4q(r^{2}+q^{2})}+\frac{\gamma_{1}Q^{2}}{4(r^{2}+q^{2})},$ (10)
where the regularization parameter $q$ corresponds to the radius of the wormhole throat and can be interpreted as the scalar charge; $Q$ is an integration constant that contributes to both the mass and electric charge of the EMP spacetime. Note that the EMP spacetime is asymptotically flat and regular, featuring a wormhole throat (also called a bounce) at $r=0$. The mass $M$ and electric charge $Q_{\mathrm{e}}$ are
$M=\frac{\gamma_{2}Q^{2}}{8q},\qquad Q_{\mathrm{e}}=\gamma_{1}Q.$ (11)
If the following condition is chosen,
$\left(\frac{Q_{\mathrm{e}}}{\gamma_{1}}\right)^{2}>\frac{8M^{2}\left(-\gamma_{1}+\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}\right)}{\gamma_{2}^{2}},$ (12)
no real zeros of $h(r)$ exist and the geometry describes a wormhole; otherwise, if
$\left(\frac{Q_{\mathrm{e}}}{\gamma_{1}}\right)^{2}\leqslant\frac{8M^{2}\left(-\gamma_{1}+\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}\right)}{\gamma_{2}^{2}},$ (13)
with the equality marking the extremal limit, the real zeros $r_{\pm}$ appear, corresponding to the horizons of a regular black hole,
$r_{\pm}=\frac{\gamma_{2}^{2}Q^{2}\pm\sqrt{\gamma_{2}^{2}Q^{4}-16\gamma_{1}q^{2}Q^{2}-64q^{4}}}{8q}.$ (14)
The horizon distribution of EMP black holes admits two distinct configurations: Type I resembles that of Reissner-Nordström (RN) black holes, where the two horizons locate on one side of the bounce; Type II possesses the horizon distribution of a bounce geometry that smoothly interpolates between the black hole and the white hole regions. The EMP framework reduces to general relativity in the classical limit,
$\gamma_{1}=\frac{1}{4},\qquad\gamma_{2}=\frac{8qM}{Q^{2}},\qquad q\to 0^{+}.$ (15)
This correspondence arises from the fundamental principles of general relativity, e.g., the preservation of U(1) gauge symmetry, enforcement of the no-scalar-hair theorem, and satisfaction of the positive mass theorem [86], thereby establishing theoretical consistency of the EMP theory with the Einstein-Maxwell dynamics.
3 Perturbation equations
3.1 General set up
In this section, we study the linear perturbation of the EMP spacetime. The phantom scalar field, Maxwell field and gravitational field as perturbations can be written as
$\phi$ $\to\bar{\phi}+\delta\phi,$ (16)
$A_{\mu}$ $\to\bar{A}_{\mu}+\delta A_{\mu},$ (17)
$g_{\mu\nu}$ $\to\bar{g}_{\mu\nu}+\delta g_{\mu\nu},$ (18)
where the quantities with a bar represent background fields, and those with a $\delta$ denote the perturbation fields. Following Chandrasekhar’s procedure [87], the perturbed spacetime is characterized by a general non-stationary axisymmetric line element,
$\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}s{{}^{2}}=-\mathrm{e}{{}^{2\nu}}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}t{{}^{2}}+\mathrm{e}{{}^{2\psi}}\left(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\varphi-q{{}_{2}}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x^{2}-q{{}_{3}}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x^{3}-\sigma\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}t\right)^{2}+\mathrm{e}{{}^{2\mu_{2}}}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x^{2})^{2}+\mathrm{e}{{}^{2\mu_{3}}}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}x^{3})^{2},$ (19)
where $q_{2}$, $q_{3}$, $\sigma$, $\nu$, $\psi$, $\mu_{2}$ and $\mu_{3}$ are functions of spacetime coordinates ($t,r,\theta$) and admit the freedom of gauge [87],
$(\sigma_{,2}-q_{2,0})_{,3}-(\sigma_{,3}-q_{3,0})_{,2}+(q_{2,3}-q_{3,2})_{,0}=0,$ (20)
where the comma denotes a partial derivative with respect to the local coordinates.
The charged Ellis wormhole solution of the EMP theory [84], as a special case of Eq. (19), has the following coefficients,
$e^{2\nu}=e^{-2\mu_{2}}=h(r)=\frac{\Delta}{\varrho^{2}},\quad e^{2\psi}=\varrho^{2}\sin^{2}{\theta},\quad e^{2\mu_{3}}=\varrho^{2},\quad q_{2}=q_{3}=\sigma=0,\quad\varrho^{2}=r^{2}+q^{2}.$ (21)
Consequently, a general perturbation of the EMP spacetime gives rise to non-vanishing $q_{2}$, $q_{3}$ and $\sigma$, which are small quantities of the linear order that introduce a rotation to the spacetime and are referred to as axial (odd-parity) gravitational field perturbations. Meanwhile, the functions $\nu$, $\mu_{2}$, $\mu_{3}$ and $\psi$ acquire small increments $\delta\nu$, $\delta\mu_{2}$, $\delta\mu_{3}$ and $\delta\psi$, which do not induce any rotation and are known as polar (even-parity) gravitational field perturbations.
In accordance with Chandrasekhar’s procedure [87], we also implement the tetrad formalism using the basis covectors,
[…]
The corresponding basis vectors can be obtained by contracting the basis covectors in Eq. (22) with the inverse of metric Eq. (19). Such a formalism determines a local inertial frame with the Minkowski metric, $\eta_{ab}={e_{a}}^{\mu}e_{b\mu}$.
To continue discussions for simplicity, we set
$P_{ab}={e_{a}}^{\mu}{e_{b}}^{\nu}Z^{-1}F_{\mu\nu},$ (23)
so that the equation of motion Eq. (6) takes the standard form,
$\nabla^{a}P_{ab}=0,$ (24)
which consists of four component equations,
$(\mathrm{e}^{\psi+\mu_{3}}P_{02})_{,2}+(\mathrm{e}^{\psi+\mu_{2}}P_{03})_{,3}$ $=0,$
$(\mathrm{e}^{\psi+\mu_{2}}P_{03})_{,0}-(\mathrm{e}^{\psi+\nu}P_{23})_{,2}$ $=0,$ (25)
$(\mathrm{e}^{\psi+\mu_{3}}P_{02})_{,0}+(\mathrm{e}^{\psi+\nu}P_{23})_{,3}$ $=0,$
$(\mathrm{e}^{\mu_{2}+\mu_{3}}P_{01})_{,0}+(\mathrm{e}^{\nu+\mu_{3}}P_{12})_{,2}+(\mathrm{e}^{\nu+\mu_{2}}P_{13})_{,3}$ $=\mathrm{e}^{\psi+\mu_{3}}P_{02}Q_{02}+\mathrm{e}^{\psi+\mu_{2}}P_{03}Q_{03}-\mathrm{e}^{\psi+\nu}P_{23}Q_{23},$
where $Q_{AB}\equiv q_{A,B}-q_{B,A}$ and $Q_{A0}\equiv q_{A,0}-\sigma_{,A}$ with $A,B=2,3$. Although $P_{ab}$ does not satisfy the Bianchi identity and is thus not the true Maxwell field, we refer to $P_{ab}$ as the electromagnetic field in our perturbation analysis and reserve the term Maxwell field exclusively for the genuine field $F_{ab}$.
The Bianchi identity for the Maxwell field,
$\nabla_{[c}F_{ab]}=0,$ (26)
also has four component equations,
$(\mathrm{e}^{\psi+\mu_{2}}F_{12})_{,3}+(\mathrm{e}^{\psi+\mu_{3}}F_{31})_{,2}$ $=0,$
$(\mathrm{e}^{\psi+\nu}F_{01})_{,2}+(\mathrm{e}^{\psi+\mu_{2}}F_{12})_{,0}$ $=0,$ (27)
$(\mathrm{e}^{\psi+\nu}F_{01})_{,3}+(\mathrm{e}^{\psi+\mu_{3}}F_{13})_{,0}$ $=0,$
$(\mathrm{e}^{\nu+\mu_{2}}F_{02})_{,3}-(\mathrm{e}^{\nu+\mu_{3}}F_{03})_{,2}+(\mathrm{e}^{\mu_{2}+\mu_{3}}F_{23})_{,0}$ $=\mathrm{e}^{\psi+\nu}F_{01}Q_{23}+\mathrm{e}^{\psi+\mu_{2}}F_{12}Q_{03}-\mathrm{e}^{\psi+\mu_{3}}F_{13}Q_{02}.$
Eqs. (25) and (27) can be reorganized into two sets of four equations, one for the axial sector and the other for the polar sector, each containing only three independent degrees of freedom.
3.2 Phantom scalar field perturbation
At first, we deal with the phantom scalar field perturbation. By linearizing the equation of motion Eq. (5), one obtains the master equation of phantom scalar field perturbations,
$\square\delta\phi$ $=\frac{1}{4}\left(\frac{\partial^{2}Z^{-1}}{{\partial\phi}^{2}}F^{2}\delta\phi+\frac{\partial Z^{-1}}{\partial\phi}\delta F^{2}\right).$ (28)
The time independence and the spherical symmetry of the EMP metric given by Eqs. (9) and (10) imply the decomposition of variables,
$\delta\phi(t,r,\theta)=\sum_{l,m}\mathrm{e}^{-\mathrm{i}\omega t}\frac{\Psi(r)}{\varrho}Y_{lm}(\theta).$ (29)
Substituting the EMP metric into Eq. (28) then yields the radial master equation,
$h^{2}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}^{2}\Psi}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r^{2}}+hh^{\prime}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\Psi}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r}+\left(\omega^{2}-V\right)\Psi=0,$ (30)
with the effective potential
$V=h\left(\frac{rh^{\prime}}{\varrho^{2}}+\frac{q^{2}h}{\varrho^{4}}+\frac{l(l+1)}{\varrho^{2}}+\frac{Q^{2}}{8q^{2}\varrho^{2}}\left(2rZ^{\prime}-\varrho^{2}Z^{\prime\prime}\right)\right),$ (31)
where a prime denotes the derivative with respect to the radial coordinate. If the last term is omitted in Eq. (31), we reduce this effective potential to the case of massless scalar field perturbations in the EMP spacetime as already presented in Ref. [78].
3.3 Gravito-Electromagnetic field perturbation
In the gravito-electromagnetic field perturbation, the linearization of the Einstein equation Eq. (7) yields
$\delta R_{ab}$ $=\frac{1}{2}\biggl{(}F_{ab}^{2}-\frac{1}{4}\eta_{ab}F^{2}\biggr{)}\frac{\partial Z^{-1}}{\partial\phi}\delta\phi+\frac{1}{2}Z^{-1}\biggl{(}\delta F_{am}{F_{b}}^{m}+F_{am}\delta{F_{b}}^{m}-\frac{1}{2}\eta_{ab}F^{mn}\delta F_{mn}\biggr{)}$
$\quad-\frac{1}{2}\delta\left(\partial_{a}\phi\partial_{b}\phi\right).$ (32)
The left-hand side of the above equation can be computed via Cartan’s equations of structure applied to the 1-forms in Eq. (22). In particular, the axial components $\delta R_{12}$ and $\delta R_{13}$ take the forms,
$(\sigma_{,2}-q_{2,0})_{,0}=-\frac{1}{\varrho^{4}\sin^{3}{\theta}}\frac{\partial\mathcal{Q}}{\partial\theta}+\frac{QZe^{\nu}}{\varrho^{3}\sin^{2}{\theta}}B,$ (33)
$(\sigma_{,3}-q_{3,0})_{,0}=\frac{\Delta}{\varrho^{4}\sin^{3}{\theta}}\frac{\partial\mathcal{Q}}{\partial r},$ (34)
where $\mathcal{Q}(r,\theta,t)=\varrho^{2}\mathrm{e}^{2\nu}Q_{23}\sin^{3}\theta$ encodes gravitational field perturbations and $B(r,\theta)=\delta P_{01}\sin{\theta}$ ) represents electromagnetic field perturbations. Note that those terms involving $\delta\left(\partial_{a}\phi\partial_{b}\phi\right)$ vanish in Eqs. (33) and (34) by virtue of the axial symmetry of the background metric Eq. (19). Eliminating $\sigma$ in Eqs. (33) and (34) and assuming a time dependence $\mathrm{e}^{\mathrm{i}\omega t}$ in the perturbations, we find the first master equation for gravito-electromagnetic field perturbations,
$\varrho^{4}\frac{\partial}{\partial r}\left(\frac{\Delta}{\varrho^{4}}\frac{\partial\mathcal{Q}}{\partial r}\right)+\sin^{3}{\theta}\frac{\partial}{\partial\theta}\left(\frac{1}{\sin^{3}{\theta}}\frac{\partial\mathcal{Q}}{\partial\theta}\right)+\omega^{2}\frac{\varrho^{4}}{\Delta}\mathcal{Q}=QZ\mathrm{e}^{\nu}\varrho\sin^{3}{\theta}\frac{\partial}{\partial\theta}\left(\frac{B}{\sin^{2}{\theta}}\right).$ (35)
As mentioned earlier, the axial electromagnetic field equations have three independent components, so as to their linearized forms,
$(\mathrm{e}^{\psi+\nu}\delta F_{01})_{,2}+(\mathrm{e}^{\psi+\mu_{2}}\delta F_{12})_{,0}=0,$
$(\mathrm{e}^{\psi+\nu}\delta F_{01})_{,3}+(\mathrm{e}^{\psi+\mu_{3}}\delta F_{13})_{,0}=0,$ (36)
$(\mathrm{e}^{\mu_{2}+\mu_{3}}\delta P_{01})_{,0}+(\mathrm{e}^{\nu+\mu_{3}}\delta P_{12})_{,2}+(\mathrm{e}^{\nu+\mu_{2}}\delta P_{13})_{,3}=\mathrm{e}^{\psi+\mu_{3}}P_{02}Q_{02}.$
Combining the above three equations and eliminating $\sigma$ yields the second master equation for gravito-electromagnetic field perturbations,
$).\displaystyle\left[Z^{-1}\mathrm{e}^{2\nu}\left(Z\varrho\mathrm{e}^{\nu}B\right)_{,2}\right]_{,2}+\frac{\mathrm{e}^{\nu}}{\varrho}\left(\frac{B_{,3}}{\sin{\theta}}\right)_{,3}\sin{\theta}+\left(\omega^{2}\varrho\mathrm{e}^{-\nu}-\frac{Q^{2}}{\varrho^{3}}Z\mathrm{e}^{\nu}\right)B=-Q\frac{\mathcal{Q}_{,3}}{\varrho^{4}\sin{\theta}}.$ (37)
The variables $r$ and $\theta$ in Eqs. (35) and (37) can be separated by the substitutions,
$\mathcal{Q}(r,\theta)=\mathcal{Q}(r)C_{l+2}^{-3/2}(\theta),$ (38)
$B(r,\theta)=\frac{B(r)}{\sin\theta}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}C_{l+2}^{-3/2}(\theta)}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\theta}=3B(r)C_{l+1}^{-1/2}(\theta),$ (39)
where the Gegenbauer function $C_{\nu}^{n}(\theta)$ satisfies
$\left[\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\theta}\sin^{2\nu}\theta\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\theta}+n(n+2\nu)\sin^{2\nu}\theta\right]C_{n}^{\nu}(\theta)=0,$ (40)
and has the recurrence relation,
$\frac{1}{\sin\theta}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}C_{n}^{\nu}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}\theta}=-2\gamma C_{n-1}^{\nu+1}.$ (41)
After separating variables, the gravito-electromagnetic field perturbation is governed by the coupled master equations,
$\Delta\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r}\left(\frac{\Delta}{\varrho^{4}}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r}\mathcal{Q}(r)\right)-\mu^{2}\frac{\Delta}{\varrho^{4}}\mathcal{Q}(r)+\omega^{2}\mathcal{Q}(r)=-\frac{QZ\mu^{2}}{\varrho^{3}}\Delta\mathrm{e}^{\nu}B(r),$ (42)
$\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r}\left[Z^{-1}\mathrm{e}^{2\nu}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r}\left(Z\varrho\mathrm{e}^{\nu}B(r)\right)\right]-(\mu^{2}+2)\frac{\mathrm{e}^{\nu}}{\varrho}B(r)+\left(\omega^{2}\varrho\mathrm{e}^{-\nu}-\frac{Q^{2}}{\varrho^{3}}Z\mathrm{e}^{\nu}\right)B(r)=-\frac{Q}{\varrho^{4}}\mathcal{Q}(r),$ (43)
where $\mu^{2}=(l-1)(l+2)$. To eliminate the first-order derivative, we introduce two functions $H_{1}$ and $H_{2}$,
$\sqrt{Z}\varrho\mathrm{e}^{\nu}B(r)=-\frac{H_{1}}{\mu},\qquad\mathcal{Q}(r)=\varrho H_{2},$ (44)
so that $H_{1}$ and $H_{2}$ represent electromagnetic and gravitational field perturbations, respectively, with subscripts denoting the spin of the fields. We further introduce $H_{0}=\Psi(r)$ and adopt the tortoise coordinate $r_{*}$ defined by
$\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r_{*}=\frac{1}{h(r)}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r,$ (45)
and then change the master equation to the Schrödinger-like form,
$\left(\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}^{2}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r_{*}^{2}}+\omega^{2}-\mathbf{V}\right)\mathbf{H}=0,$ (46)
where
$\mathbf{H}=\left(\begin{array}[]{c}H_{0}\\ H_{1}\\ H_{2}\end{array}\right),\qquad\mathbf{V}=\left(\begin{array}[]{ccc}V&0&0\\ 0&V_{11}&V_{12}\\ 0&V_{21}&V_{22}\end{array}\right),$ (53)
and
$V_{11}$ $=\frac{3}{4Z^{2}}\left(\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}Z}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r_{*}}\right)^{2}-\frac{1}{2Z}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}^{2}Z}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r_{*}^{2}}+\left(\mu^{2}+2\right)\frac{\Delta}{\varrho^{4}}+\frac{ZQ^{2}\Delta}{\varrho^{6}},$
$V_{12}$ $=V_{21}=\frac{Q\sqrt{Z}\mu\Delta}{\varrho^{5}},$
$V_{22}$ $=\mu^{2}\frac{\Delta}{\varrho^{4}}+\frac{2\Delta^{2}r^{2}}{\varrho^{8}}-\frac{\Delta}{\varrho^{3}}\frac{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}}{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}r}\left(\frac{\Delta r}{\varrho^{3}}\right).$ (54)
In the Reissner-Nordström (RN) background, the effective potential matrix in the master equation can be diagonalized by a constant matrix, yielding decoupled perturbation equations. However, the coupling cannot be removed in this manner [^1] It has come to our attention that some studies [88, 89, 90] have inappropriately applied the RN decoupling procedure in their work. To avoid further misleading, we explicitly clarify here that the RN decoupling procedure is not valid in the EMP spacetime. in the EMP spacetime, so we shall apply numerical methods directly to the coupled system.
4 Instability of wormholes
4.1 Singularity of the gravito-electromagnetic effective potential
As we have mentioned before, the coupling function $Z$ is not positive definite and vanishes at
$\zeta_{\pm}=$ $-\frac{q}{\gamma_{1}}\left(\gamma_{2}\pm\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}\right),$ (55)
introducing the divergent behavior in $V_{11}$. In the EMP spacetime, $\zeta_{\pm}$ mark critical points at which the electromagnetic field transitions from a conventional regime on one side to a phantom-like regime on the other side and can cause the effective potential $V_{11}$ to diverge due to the presence of $Z$ on the denominator. In the Type I black hole scenario, these divergences are strictly confined within the Cauchy horizon ($r_{-}>\zeta_{+}$), a condition equivalent to
$8\gamma_{2}q^{2}+\gamma_{1}\gamma_{2}Q^{2}>\gamma_{1}\sqrt{\gamma_{2}^{2}Q^{4}-16\gamma_{1}q^{2}Q^{2}-64q^{4}}+8q^{2}\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}.$ (56)
Through two consecutive squaring operations applied to both sides, the inequality Eq. (56) transforms into the manifestly satisfied condition,
$\left(\gamma_{1}^{2}+\gamma_{2}^{2}\right)\left(8q^{2}+\gamma_{1}Q^{2}\right)^{2}>0.$ (57)
Hence, the divergence points of $V_{11}$ are proven to be hidden behind the Cauchy horizon, which indicates the stability of EMP black holes. Nevertheless, as the electric charge increases beyond the extremal value, the black hole undergoes a transition to a wormhole configuration, where singularities of $V_{11}$ become exposed to the outer spacetime, violating cosmic censorship and rendering the configuration physically inadmissible.
4.2 Finite difference method
The finite difference method is widely used to simulate the time evolution of perturbations [91, 77]. For this purpose, we grid the spacetime coordinates $t=i\Delta t$ and $r_{*}=j\Delta r_{*}$, where $i$ and $j$ are both integers. Taking the step length $\Delta t$ and $\Delta r_{*}$ to be constant, $\mathbf{H}\left(t,r_{*}\right)$ and $\mathbf{V}\left(t,r_{*}\right)$ can be abbreviated as $\mathbf{H}\left(i,j\right)$ and $\mathbf{V}\left(i,j\right)$. The master equation Eq. (46) can therefore be expressed as
$-\frac{\mathbf{H}\left(i+1,j\right)-2\mathbf{H}\left(i,j\right)+\mathbf{H}\left(i-1,j\right)}{\Delta t^{2}}+\frac{\mathbf{H}\left(i,j+1\right)-2\mathbf{H}\left(i,j\right)+\mathbf{H}\left(i,j-1\right)}{\Delta r_{*}^{2}}$
$-\mathbf{V}\left(j\right)\mathbf{H}\left(i,j\right)=0.$ (58)
We consider the initial Gaussian distribution,
$\mathbf{H}\left(t=0,r_{*}\right)=\left(\begin{array}[]{c}\mathrm{e}^{-\frac{\left(r_{*}-a\right)^{2}}{2b^{2}}}\\ \mathrm{e}^{-\frac{\left(r_{*}-a\right)^{2}}{2b^{2}}}\\ 0\end{array}\right),\qquad\mathbf{H}\left(t<0,r_{*}\right)=\mathbf{0},$ (62)
where $a$ and $b$ are the median and width of the initial wave packet, and then calculate the time evolution of the perturbation via the iteration scheme,
$\mathbf{H}\left(i+1,j\right)=-\mathbf{H}\left(i-1,j\right)+\left[2-2\frac{\Delta t^{2}}{\Delta r_{*}^{2}}-\Delta t^{2}\mathbf{V}\left(j\right)\right]\mathbf{H}\left(i,j\right)+\frac{\Delta t^{2}}{\Delta r_{*}^{2}}\left[\mathbf{H}\left(i,j+1\right)+\mathbf{H}\left(i,j-1\right)\right].$ (63)
To satisfy the von Neumann stability condition, $\Delta t/\Delta r_{*}<1$, we set $\Delta t=0.1$, $\Delta r_{*}=0.2$, $a=20$, and $b=1$.
4.3 Divergence of phantom scalar field perturbations
A previous study claimed [78] that EMP wormholes remain stable under massless scalar field perturbations. However, the phantom scalar field, serving as the matter source in the EMP spacetime, carries negative kinetic energy and may therefore induce spacetime instabilities.
Fig. 1 shows how the azimuthal number $l$ affects the effective potential and time domain waveform for phantom scalar field perturbations. For $l=0$, the effective potential contains a negative region, causing the waveform to diverge rapidly after the initial ringing. As $l$ increases, this region disappears and the effective potential develops a double-peak structure, yielding echo signals in the waveform. These echoes follow the patterns remarkably similar to those of a minimally coupled massless scalar field in Ref. [78], and will not be further discussed here. Instead, we focus on the divergence of phantom scalar field perturbations and the singular behavior of $V_{11}$ as mentioned above, both of which indicate the instability of the EMP wormhole. Here we proceed with the following possible explanations:
  • As the electric charge increases, Type I black holes go to extremal and turn into wormholes. However, extremal black holes are unattainable in general relativity, implying that neither extremal Type I black holes nor wormholes in the EMP theory are likely to be physically realized.
  • The divergence of $V_{11}$ can be resolved if the Maxwell field vanishes at $r=\zeta_{\pm}$.
  • EMP wormholes are unstable under gravitational field perturbations because of the coupling between the phantom scalar field and the Maxwell field.
  • The divergence of $V_{11}$ may arise from limitations of linearization procedure or from an inadequate choice of perturbation variables, rather than an essential singularity.
(a)  Effective potential.
(b)  Time profile.
Figure 1: The effect of the azimuthal number $l$ of EMP wormholes on the effective potential and the time profile of the phantom scalar field perturbation with $q=1.6$, $Q=1$, $\gamma_{1}=-10$ and $\gamma_{2}=2$.
As a summary to this section, we have revealed the instability of EMP wormholes from two aspects, where one aspect is the dramatic increment of $l=0$ phantom scalar field perturbations and the other is the divergence of the gravito-electromagnetic effective potential — while the former can be resolved by suitably tuning spacetime parameters, the latter persists. It is the presence of the phantom scalar field and its coupling with the Maxwell field responsible for the instability of EMP wormholes. During the wormhole-black hole transition, the event horizon forms in such a way that it hides these singularities behind the Cauchy horizon. Consequently, although the theory admits wormhole solutions, only black hole configurations with a complete horizon structure are physically viable and stable. A more detailed investigation of the underlying mechanisms and potential remedies remains an open question at present.
5 Quasinormal modes of Type I black holes
In the previous section, the instability of wormholes as demonstrated indirectly supports the stability of black holes, where the latter will be explored in great detail in this section. We will solve the master equation Eq. (46) numerically to acquire QNFs of Type I black holes. The boundary conditions for the QNM problem are pure ingoing waves at the event horizon,
$\mathbf{H}\sim\mathrm{e}^{-\mathrm{i}\omega r_{*}},\qquad r_{*}\to-\infty,$ (64)
and pure outgoing waves at the spatial infinity,
$\mathbf{H}\sim\mathrm{e}^{\mathrm{i}\omega r_{*}},\qquad r_{*}\to\infty.$ (65)
[…]
$\frac{Q_{\mathrm{e}}}{M}=\frac{1}{Q}.$ (66)
After assigning the specific electric charge, we can then vary $q$ over the interval $(0,q_{\mathrm{ext}}]$, where $q_{\mathrm{ext}}$ denotes the extremal value,
$q_{\mathrm{ext}}=\frac{Q}{4}\sqrt{Q^{2}-1}=\frac{M}{4Q_{\mathrm{e}}}\sqrt{\left(\frac{M}{Q_{\mathrm{e}}}\right)^{2}-1}.$ (67)
The QNFs are determined through two complementary methodologies: the extraction from the finite difference time domain profiles via the Prony method and the computation via a matrix-valued direct integration approach; see Refs. [57, 92] for the Prony method and Ref. [93] for the direct integration method. We represent the fundamental QNFs of the phantom scalar field, the electromagnetic field and the gravitational field perturbations of Type I black holes with $Q_{\mathrm{e}}/M=2/5$ and $q_{\mathrm{ext}}=\frac{5\sqrt{21}}{16}$ in Table 1. The azimuthal number is set to be $l=2$, corresponding to the lowest nontrivial gravitational quadrupole mode.
$q$ Phantom scalar field perturbation Electromagnetic field perturbation Gravitational field perturbation
Direct integration Prony Direct integration Prony Direct integration Prony
$\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$
0.001 0.38798  -0.089820 0.38800  -0.089832 0.47993  -0.096442 0.47994  -0.096469 0.37844  -0.089398 0.37847  -0.089422
0.2 0.38900  -0.089627 0.38902  -0.089635 0.48187  -0.096421 0.48192  -0.096457 0.37810  -0.089208 0.37813  -0.089232
0.4 0.39213  -0.089033 0.39214  -0.089042 0.48821  -0.096495 0.48828  -0.096525 0.37725  -0.088654 0.37728  -0.088687
0.6 0.39759  -0.087988 0.39760  -0.087999 0.49874  -0.096263 0.49879  -0.096305 0.37614  -0.087714 0.37617  -0.087736
0.8 0.40583  -0.086378 0.40584  -0.086388 0.51290  -0.095261 0.51297  -0.095300 0.37497  -0.086303 0.37500  -0.086323
1 0.41765  -0.083942 0.41768  -0.083955 0.53111  -0.093153 0.53113  -0.093161 0.37391  -0.084267 0.37392  -0.084265
1.2 0.43460  -0.079995 0.43461  -0.080005 0.55485  -0.089156 0.55493  -0.089197 0.37302  -0.081359 0.37304  -0.081379
1.4 0.45982  -0.072113 0.45984  -0.072124 0.58734  -0.080641 0.58743  -0.080682 0.37208  -0.077231 0.37210  -0.077250
1.43 0.46468  -0.070129 0.46466  -0.070136 0.59332  -0.078506 0.59339  -0.078546 0.37190  -0.076496 0.37192  -0.076515
Table 1: Fundamental QNMs of the phantom scalar field perturbation, the electromagnetic field perturbation and the gravitational field perturbation for Type I black holes with parameters $\gamma_{1}=1/4$, $\gamma_{2}=2q$ and $Q_{\mathrm{e}}/M=2/5$, computed via the direct integration method and the Prony method for a varying scalar charge $q$. The azimuthal number $l$ is fixed at $l=2$.
A critical consistency check emerges in the vanishing scalar charge limit, where both the master equations Eqs. (42) and (43) as well as the corresponding QNFs reproduce the case of axial (odd-parity) gravito-electromagnetic field perturbations of RN black holes. This concordance (notably, the cases of three field perturbations at $q=0.001$ in Table 1 match the RN predictions) provides a robust validation of our calculation. Moreover, the agreement between the direct integration and Prony results confirms the validity of the numerical methods. Although the direct integration method is highly efficient for simple perturbation equations like the phantom scalar field perturbation equation, see Eq. (30), it becomes time-consuming for complex ones like the gravito-electromagnetic coupled perturbation equations, see Eqs. (42) and (43). Accordingly, we employ the direct integration method for phantom scalar field perturbations and use the finite difference method combined with the Prony method for gravito-electromagnetic field perturbations.
Fig. 2 demonstrates that the scalar charge $q$ induces distinct spectral shifts in QNFs for different field perturbations.
  • For the phantom scalar and electromagnetic field perturbations, the real frequency $\omega_{\mathrm{R}}$ increases when $q$ grows, whereas it decreases slightly for the gravitational field perturbation.
  • The imaginary part $\omega_{\mathrm{I}}$ of QNFs grows monotonically for both the phantom scalar and gravitational field perturbations as $q$ rises, suggesting that a larger scalar charge leads to a longer lived ringdown signal. By contrast, it exhibits a non-monotonic dependence for the electromagnetic field perturbation: $\omega_{\mathrm{I}}$ initially decreases at a small scalar charge $q$ and then increases at a large $q$.
(a)  Real parts of QNFs.
(b)  Imaginary parts of QNFs.
Figure 2: The effect of the scalar charge $q$ of Type I black holes on the fundamental QNFs of the phantom scalar field perturbation with $\gamma_{1}=1/4$, $\gamma_{2}=2q$, $Q_{\mathrm{e}}/M=2/5$ and $l=2$.
Next we investigate how the relative contributions of the phantom scalar and electromagnetic field perturbations affect the QNMs. By rescaling $q$ and $Q_{\mathrm{e}}$ by the mass $M$, i.e., $\tilde{q}=q/M$ and $\tilde{Q}_{\mathrm{e}}=Q_{\mathrm{e}}/M$, we further define the generalized specific charge,
$\eta=\frac{\sqrt{q^{2}+Q_{\mathrm{e}}^{2}}}{M}=\sqrt{\tilde{q}^{2}+\tilde{Q}_{\mathrm{e}}^{2}},$ (68)
where $\eta=1$ corresponds to extremal black holes. We introduce a mixing angle between the scalar and electric charges by
$\tan{\vartheta}\equiv\frac{{q}}{{Q}_{\mathrm{e}}}=\frac{\tilde{q}}{\tilde{Q}_{\mathrm{e}}},$ (69)
where $\vartheta\in[0,\pi/2]$, and express the dimensionless charges using $\eta$ and $\vartheta$,
$\tilde{q}=\eta\sin{\vartheta},\qquad\tilde{Q}_{\mathrm{e}}=\eta\cos{\vartheta}.$ ) . (70)
If $\eta$ is fixed but $\vartheta$ varied, we can keep the horizon radius unchanged but adjust the relative contributions of the scalar and electric charges.
Figs. 3-5 illustrate how the fundamental QNFs depend on the charge composition for the phantom scalar field, electromagnetic field, and gravitational field perturbations of Type I black holes, respectively. The real frequency and damping rate of the perturbed phantom scalar and gravitational fields, see Figs. 3 and 5, decrease monotonically as the mixing angle $\vartheta$ increases, in particular, the amplitudes become smaller for larger values of the generalized specific charge $\eta$. For the perturbed electromagnetic field, see Fig. 4, the real frequency and damping rate show a different tendency: $\omega_{\mathrm{R}}$ increases monotonically as the mixing angle $\vartheta$ increases at a low $\eta$ but decreases monotonically at a high $\eta$; the damping rate rises when the mixing angle is small at a low $\eta$, but it decreases monotonically when the mixing angle increases at a high $\eta$. The values of QNFs are given in Tables 2-4.
(a)  Real parts of QNFs.
(b)  Imaginary parts of QNFs.
Figure 3: The effect of the mixing angle $\vartheta$ of Type I black holes on the fundamental QNFs of the phantom scalar field perturbation with $l=2$.
(a)  Real parts of QNFs.
(b)  Imaginary parts of QNFs.
Figure 4: The effect of the mixing angle $\vartheta$ of Type I black holes on the fundamental QNFs of the electromagnetic field perturbation with $l=2$.
(a)  Real parts of QNFs.
(b)  Imaginary parts of QNFs.
Figure 5: The effect of the mixing angle $\vartheta$ of Type I black holes on the fundamental QNFs of the gravitational field perturbation with $l=2$.
$\vartheta$ $\eta$ $0.1$ $0.3$ $0.5$ $0.7$ $0.9$
$\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$
0.1 0.37450   -0.089016 0.38143   -0.089438 0.39697   -0.090222 0.42607   -0.091038 0.48534   -0.089374
0.3 0.37448   -0.089004 0.38121   -0.089324 0.39624   -0.089885 0.42429   -0.090306 0.48033   -0.088016
0.5 0.37444   -0.088981 0.38079   -0.089116 0.39492   -0.089275 0.42091   -0.089005 0.47155   -0.085702
0.7 0.37438   -0.088952 0.38025   -0.088847 0.39322   -0.088501 0.41671   -0.087404 0.46091   -0.083011
0.9 0.37432   -0.088921 0.37968   -0.088562 0.39142   -0.087694 0.41236   -0.085787 0.45036   -0.080450
1.1 0.37427   -0.088892 0.37916   -0.088305 0.38981   -0.086981 0.40853   -0.084399 0.44144   -0.078364
1.3 0.37423   -0.088871 0.37877   -0.088116 0.38863   -0.086462 0.40575   -0.083415 0.43519   -0.076940
1.5 0.37421   -0.088861 0.37858   -0.088023 0.38804   -0.086209 0.40440   -0.082941 0.43220   -0.076270
Table 2: Fundamental QNMs of the phantom scalar field perturbation for Type I black holes with $\gamma_{1}=1/4$, where the direct integration method is used for different values of $\vartheta$ and $\eta$. The azimuthal number $l$ is set to be $l=2$.
$\vartheta$ $\eta$ $0.1$ $0.3$ $0.5$ $0.7$ $0.9$
$\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$
0.1 0.45898   -0.095134 0.46989   -0.095854 0.49364   -0.097200 0.53630   -0.098740 0.61865   -0.097479
0.3 0.45906   -0.095178 0.46978   -0.095819 0.49293   -0.096978 0.53413   -0.098152 0.61209   -0.096264
0.5 0.45964   -0.095427 0.46994   -0.095908 0.49199   -0.096700 0.53049   -0.097197 0.60083   -0.094229
0.7 0.46103   -0.095904 0.47078   -0.096157 0.49132   -0.096449 0.52634   -0.096069 0.58754   -0.091852
0.9 0.46319   -0.096360 0.47225   -0.096361 0.49110   -0.096113 0.52240   -0.094838 0.57467   -0.089498
1.1 0.46553   -0.096538 0.47394   -0.096320 0.49117   -0.095609 0.51912   -0.093601 0.56397   -0.087451
1.3 0.46741   -0.096458 0.47530   -0.096093 0.49130   -0.095073 0.51680   -0.092586 0.55655   -0.085962
1.5 0.46836   -0.096345 0.47599   -0.095914 0.49137   -0.094753 0.51567   -0.092051 0.55301   -0.085230
Table 3: Fundamental QNMs of the electromagnetic field perturbation for Type I black holes with $\gamma_{1}=1/4$, where the Prony method is used for different values of $\vartheta$ and $\eta$. The azimuthal number $l$ is set to be $l=2$.
$\vartheta$ $\eta$ $0.1$ $0.3$ $0.5$ $0.7$ $0.9$
$\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$       $\omega_{\mathrm{I}}M$ $\omega_{\mathrm{R}}M$      $\omega_{\mathrm{I}}M$
0.1 0.37396   -0.089014 0.37618   -0.089226 0.38157   -0.089602 0.39223   -0.089855 0.41297   -0.088231
0.3 0.37392   -0.089004 0.37581   -0.089129 0.38049   -0.089322 0.38990   -0.089269 0.40813   -0.087251
0.5 0.37384   -0.088985 0.37514   -0.088959 0.37858   -0.088830 0.38580   -0.088251 0.39982   -0.085599
0.7 0.37375   -0.088962 0.37432   -0.088749 0.37623   -0.088225 0.38083   -0.087021 0.39008   -0.083693
0.9 0.37366   -0.088938 0.37349   -0.088527 0.37387   -0.087598 0.37588   -0.085784 0.38075   -0.081874
1.1 0.37358   -0.088915 0.37278   -0.088320 0.37184   -0.087027 0.37171   -0.084699 0.37312   -0.080355
1.3 0.37353   -0.088897 0.37227   -0.088158 0.37040   -0.086591 0.36878   -0.083899 0.36791   -0.079279
1.5 0.37350   -0.088887 0.37202   -0.088074 0.36971   -0.086369 0.36737   -0.083501 0.36545   -0.078755
Table 4: Fundamental QNMs of the gravitational field perturbation for Type I black holes with $\gamma_{1}=1/4$, where the Prony method is used for different values of $\vartheta$ and $\eta$. The azimuthal number $l$ is set to be $l=2$.
6 Conclusion
The EMP theory can be derived from the EMD theory, where the latter is a four-dimensional low-energy effective theory of string theory. In the EMP spacetime, the phantom scalar field is minimally coupled with Einstein gravity but nonminimally coupled with the Maxwell field. A static and spherically symmetric solution, the so-called charged Ellis wormhole also describes a regular black hole, providing us a frame to study wormholes or black holes under this spacetime background.
We investigate the QNMs of the EMP spacetime by studying the linear perturbation of the metric following Chandrasekhar’s procedure. Because of the spherical symmetry of the EMP spacetime, the radial parts of the perturbed fields can be decomposed from the angular parts, and the latter can be expanded by the spherical harmonics. The perturbation can be classified into the axial (odd-parity) sector and polar (even-parity) sector based on their dependence on the angular coordinate $\varphi$. In the axial sector, the gravitational field perturbation and the electromagnetic field perturbation are coupled, whereas the phantom scalar field perturbation is decoupled. We derive the corresponding three master equations and rewrite the resulting Schrödinger-like equations in a compact matrix form, see Eq. (46). We notice that both the Lagrangian and metric of the EMP theory reduce to general relativity when the phantom scalar hair vanishes and the coupling parameters are chosen properly, so do the master equations and QNFs.
Because of the coupling between the Maxwell field and the phantom scalar field, the effective potential $V_{11}$ contains the coupling function $Z$ in the denominator of $V_{11}$, see Eq. (54), causing $V_{11}$ to diverge at the zeros of $Z$. We show that these divergence points lie strictly inside the Cauchy horizon in the black hole scenario, whereas they are exposed to the exterior region in the wormhole case. Moreover, the effective potential contains a negative region when the EMP wormhole is perturbed by a phantom scalar field in the case of $l=0$, leading to a growing time profile obtained via the finite difference method, see Fig. 1. These features not only manifest the instability of EMP wormholes but also, by contrast, underscore the perturbative stability of EMP black holes.
We extract the fundamental QNFs from finite difference time profiles using the Prony method and also compute them via the matrix-valued direct integration method. Specifically, We choose $\gamma_{1}=1/4$ and $\gamma_{2}=2q$, so that the Type I black hole mimics the RN black hole as the scalar charge $q\to 0^{+}$. Table 1 tabulates the values of the fundamental QNFs under the phantom scalar field perturbation, the electromagnetic field perturbation and the gravitational field perturbation, respectively, for different values of the scalar charge $q$ with a fixed specific electric charge $\tilde{Q}_{\mathrm{e}}=2/5$, and Fig. 2 provides the corresponding graphical depiction. The results obtained from the Prony method and the matrix-valued direct integration method agree well with each other, which strengthens the validity of our results. As expected, the QNFs are consistent with those of the RN black hole when $q=0.001$.
By introducing the generalized specific charge $\eta$ and the mixing angle $\vartheta$, we are able to study the effect of relative contributions of the scalar and electric charges on the QNMs of Type I black holes, see Figs. 3-5 and Tables 2-4. We find that the real frequencies and damping rates behave similarly in the phantom scalar perturbation and the gravitational field perturbation, see Figs. 3 and 5 and Tables 2 and 4 for the details, while they show a different behavior only for a small $\vartheta$ and a low $\eta$ in the electromagnetic field perturbation, see Fig. 4 and Table 3 for the details.
We focus currently on the axial perturbations of the EMP spacetime, where the interesting features on wormholes and black holes have been summarized above. We shall leave the polar perturbations in the EMP spacetime to our future studies and expect additional physical insights.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 12175108.